# Error-Minimizing Measurements in Postselected One-Shot Symmetric Quantum State Discrimination and Acceptance as a Performance Metric

### v1

@@ 3. $\underline{\mathcal{C}\mathrm{3}}$
If $\Pi_{\sigma}<\Pi_{\rho},\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{1}_{s}(\rho,\sigma)$. So, the set of all error minimizing measurement is given only by first set i.e. $\mathcal{E}_{s}^{1}(\rho,\sigma)$. Further, for the postselected symmetric error minimizing measurements, acceptance for the state $\sigma$, $A_{\sigma}(\Lambda)=0$, so if the input state is $\sigma$, the error-minimizing measurement always rejects both the hypotheses. Also, any error minimizing measurement never declares the unknown state as $\sigma$. See Fig. 5(a) for clarity.
@@
If $\Pi_{\rho}<\Pi_{\sigma},\mathcal{E}_{s}(\rho,\sigma)=\mathcal{E}^{2}_{s}(\rho,\sigma)$ and $A_{\rho}(\Lambda)=0$. So, the set of all error minimizing measurement is given only by the second set i.e. $\mathcal{E}_{s}^{2}(\rho,\sigma)$. Further, for the postselected symmetric error minimizing measurements, acceptance for the state $\rho$, $A_{\rho}(\Lambda)=0$, so if the input state is $\rho$, the error-minimizing measurement always rejects both the hypotheses. Also, any error minimizing measurement never declares the unknown state as $\rho$. See Fig. 5(b) for clarity.
@@
Starting from the definition of $\mathcal{E}_{s}^{1}(\rho,\sigma)$ as given in Theorem IV.1, we get
$\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$
$\Leftrightarrow\displaystyle\Lambda_{\rho}=\frac{c_{1}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},c_{1}>0,\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\rho}\leq\mathrm{I},\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho},\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma})$
$\Leftrightarrow\displaystyle\Lambda_{\rho}=\frac{c_{1}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),0<c_{1}\leq\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}^{-1},\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho},\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}).$
Writing it as a set, we get the stated result in the Lemma. ∎
Lemma IV.2.
@@
Starting from the definition of $\mathcal{E}_{s}^{2}(\rho,\sigma)$ as given in Theorem IV.1, we get
$\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$
$\Leftrightarrow\displaystyle\Lambda_{\sigma}=\frac{c_{2}\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)},c_{2}>0,\psi_{\sigma}\in\mathcal{S}(\mathrm{I}-\Pi_{\rho}),\Lambda_{\sigma}\leq\mathrm{I},\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma},\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma})$
$\Leftrightarrow\displaystyle\Lambda_{\sigma}=\frac{c_{2}\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)},\psi_{\sigma}\in\mathcal{S}(\mathrm{I}-\Pi_{\rho}),0<c_{2}\leq\left\|\frac{\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}^{-1},\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma},\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}).$
Writing it as a set, we get the stated result in the Lemma. ∎
Lemma IV.3.
@@
Starting from the definition of $\mathcal{E}_{s}^{3}(\rho,\sigma)$ as given in Theorem IV.1, we get
@@
Now substituting $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ in the condition $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$, we get this condition as a bound on $c_{3}$ as
@@
Writing it as a set, we get the stated result in the Lemma. ∎
@@
Note that all the three cases, taking $\psi_{\rho}$ and $\psi_{\sigma}$ such that $\mathrm{Tr}\left(\psi_{\rho}\rho\right)=0$ or $\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)=0$ is not valid, as clear from the expression of $\Lambda_{\rho}$ and $\Lambda_{\sigma}$. This ensures that $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0$ or $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0$. This is needed because $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0$ or $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0$ is among the conditions on the sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$ in Theorem IV.1.
Constructing an arbitrary error minimizing measurement: In Theorem IV.1, we saw that an arbitrary error-minimizing measurement $\Lambda$ must belong to one of $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ or $\mathcal{E}_{s}^{3}(\rho,\sigma)$. The outcome specific properties of these sets were observed in Remark 8 and illustrated in Fig. 5. Further, the sets were written in a parameterized form in the Lemma IV.1, Lemma IV.2 and Lemma IV.3. Based on these parameters, a method to construct an arbitrary error minimizing measurement is given in the table below. The first column in the table states the free variables/parameters to choose and rest give the condition on these. The conditions for $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$ are written in second, third and fourth columns respectively. To construct an arbitrary error-minimizing measurement from $\mathcal{E}_{s}^{1}(\rho,\sigma)$, start from the first row, pick an arbitrary free variable satisfying the conditions stated the second column and proceed towards the last row in this way. After getting $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ in the last row, the error-minimizing measurement is given by $\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$. For constructing measurements from the sets $\mathcal{E}_{s}^{2}(\rho,\sigma)$ or $\mathcal{E}_{s}^{3}(\rho,\sigma)$, do the same but take the conditions from the corresponding columns.
@@
TABLE II: The table shows a way to choose parameters for an arbitrary error-minimizing measurement for each of the three sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$.
Now, we give the acceptance for an arbitrary error minimizing measurement from the three sets in terms of the parameters given in Table II.
Theorem IV.2.
For an arbitrary error minimizing measurement $\Lambda\in\mathcal{E}_{s}(\rho,\sigma)$, acceptance for the states $\rho$ and $\sigma$ is given as
@@
Here $c_{1},c_{2},c_{3}$ and $c_{r}$ are parameters as given in the Table II.
@@
$\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\rho})\Rightarrow\mathrm{Tr}\left(\Gamma\rho\right)=0\text{ and }\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma})\Rightarrow\mathrm{Tr}\left(\Gamma\sigma\right)=0.$
[…]
For these three sets, the acceptance is calculated as given below.
@@
Now, we maximize the acceptance over the the set $\mathcal{E}\left(\rho,\sigma,p\right)$. To do so, we maximize the expression obtained in Theorem IV.2 with respect the parameters. The maximum acceptance expression is derived in the following theorem.
Theorem IV.3 (Maximum acceptance).
Given a pair of states as $\rho$ and $\sigma$ such that $\Pi_{\rho}\neq\Pi_{\sigma}$, then maximum acceptance for the states $\rho$ and $\sigma$ is given as
@@
$\Pi_{\rho}>\Pi_{\sigma}$ $1-\mathrm{Tr}\left(\Pi_{\sigma}\rho\right)$ 00
$\Pi_{\rho}<\Pi_{\sigma}$ 00 $1-\mathrm{Tr}\left(\Pi_{\rho}\sigma\right)$
@@
Recall from Corollary 3 that when $\Pi_{\rho}>\Pi_{\sigma}$, the set $\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{1}_{s}(\rho,\sigma)$. Similarly, from Corollary 4, we know that when $\Pi_{\rho}<\Pi_{\sigma}$, the set $\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{2}_{s}(\rho,\sigma)$. We begin with the third case i.e. $\Pi_{\rho}\not<\Pi_{\sigma}$ and $\Pi_{\rho}\not>\Pi_{\sigma}$. In this case $\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{1}_{s}(\rho,\sigma)\cup\mathcal{E}^{2}_{s}(\rho,\sigma)\cup\mathcal{E}^{3}_{s}(\rho,\sigma)$. Starting from maximizing acceptance for the state $\rho$ over the set $\mathcal{E}_{s}(\rho,\sigma,p)$, which can be obtained by maximizing over all the three sets followed by taking the maximum of the three, we get
@@
Note that $A_{\rho}(\Lambda)=0\ \forall\ \Lambda\in\mathcal{E}_{s}^{2}(\rho,\sigma)$, hence
@@
Using the expression of acceptance obtained in Theorem IV.2, we get
@@
Applying bounds on $c_{1}$ and $c_{3}$ in the set $\mathcal{E}_{s}^{1}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$, we get
@@
Note that $\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\leq\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{c_{r}\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\Rightarrow\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}\leq\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{c_{r}\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}$. Thus, the maximum is obtained by maximizing the first term and so
@@
The equality can be proved similar to proof of (75) in proof of Lemma D.2 of Appendix D. Note that among the all $\tilde{\psi}_{\rho}\in\mathcal{P}(\Pi_{\rho+\sigma}-\Pi_{\sigma})$, such that $\|\tilde{\psi}_{\rho}\|_{\infty}=1$, the one that maximizes $\mathrm{Tr}\left(\tilde{\psi}_{\rho}\rho\right)$ is $\tilde{\psi}_{\rho}=(\Pi_{\rho+\sigma}-\Pi_{\sigma})$. Hence, we obtain maximum at $\psi_{\rho}=\frac{\Pi_{\rho+\sigma}-\Pi_{\sigma}}{\mathrm{Tr}\left(\Pi_{\rho+\sigma}-\Pi_{\sigma}\right)}$ and is given by
@@
Acceptance for the state $\sigma$ is maximized in a similar way to obtain.
$A_{\sigma}^{s}$ $=1-\mathrm{Tr}\left(\Pi_{\rho}\sigma\right).$ (36)
For the case $\Pi_{\rho}<\Pi_{\sigma}$ or $\Pi_{\rho}>\Pi_{\sigma}$, the maximum acceptance can be obtained in a similar way by maximizing over the sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$ or $\mathcal{E}_{s}^{2}(\rho,\sigma)$ respectively, which make the problem simpler and we directly get expression as in (34), thus obtaining the stated values in the table. ∎
@@
In all the three cases, the stated maximum acceptance for the state $\rho$, i.e. $A_{\rho}^{s}$ is achieved by taking the measurement $\{\Pi_{\rho+\sigma}-\Pi_{\sigma},0,\mathrm{I}-\Pi_{\sigma}+\Pi_{\rho+\sigma}\}$. In all the three cases, the stated maximum acceptance for the state $\sigma$, i.e. $A_{\sigma}^{s}$ is achieved by taking the measurement $\{0,\Pi_{\rho+\sigma}-\Pi_{\rho},\mathrm{I}-\Pi_{\rho}+\Pi_{\rho+\sigma}\}$.
V Conclusion

### v2

@@ 3. $\underline{\mathcal{C}\mathrm{3}}$
If $\Pi_{\sigma}<\Pi_{\rho},\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{1}_{s}(\rho,\sigma)$. So, the set of all error-minimizing measurements is given only by the first set, i.e. $\mathcal{E}_{s}^{1}(\rho,\sigma)$. Further, for the postselected symmetric error minimizing measurements, acceptance for the state $\sigma$, $A_{\sigma}(\Lambda)=0$, so if the input state is $\sigma$, an error-minimizing measurement always rejects the test. Also, any error-minimizing measurement never declares the unknown state as $\sigma$. See Fig. 5(a) for clarity.
@@
If $\Pi_{\rho}<\Pi_{\sigma},\mathcal{E}_{s}(\rho,\sigma)=\mathcal{E}^{2}_{s}(\rho,\sigma)$ and $A_{\rho}(\Lambda)=0$. So, the set of all error-minimizing measurements is given only by the second set, i.e., $\mathcal{E}_{s}^{2}(\rho,\sigma)$. Further, for the postselected symmetric error minimizing measurements, acceptance for the state $\rho$, $A_{\rho}(\Lambda)=0$, so if the input state is $\rho$, the error-minimizing measurement always rejects both the hypotheses. Also, any error-minimizing measurement never declares the unknown state as $\rho$. See Fig. 5(b) for clarity.
@@
From the definition of $\mathcal{E}_{s}^{1}(\rho,\sigma)$ as given in Theorem 5, we get
$\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}.$
Taking $\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),c_{1}>0$, the above mentioned conditions can be parameterized as
$\displaystyle\Lambda_{\rho}=\frac{c_{1}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},c_{1}>0,\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\rho}\leq\mathrm{I},\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho},\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}).$
Using $\Lambda_{\rho}\leq\mathrm{I}$, we get an upper bound on $c_{1}$ and thus equivalent parameterization becomes
$\displaystyle\Lambda_{\rho}=\frac{c_{1}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),0<c_{1}\leq\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}^{-1},\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho},\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}).$
This gives the desired result. ∎
Lemma 2.
@@
The proof follows similar to the proof of Lemma 1. ∎
Lemma 3.
@@
From the definition of $\mathcal{E}_{s}^{3}(\rho,\sigma)$ as given in Theorem 5, we get
@@
Now, substituting $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ in the condition $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$, we get this condition as a bound on $c_{3}$ as
@@
This gives the desired result. ∎
@@
[…]
Remark 13.
Note that all in the three sets, i.e., $\mathcal{E}^{1}_{s}(\rho,\sigma),\mathcal{E}^{2}_{s}(\rho,\sigma)$ and $\mathcal{E}^{3}_{s}(\rho,\sigma)$, while parameterize measurement in terms of $\psi_{\rho}$ and $\psi_{\sigma}$, taking $\psi_{\rho}$ and $\psi_{\sigma}$ such that $\mathrm{Tr}\left(\psi_{\rho}\rho\right)=0$ or $\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)=0$ is not valid, as clear from the expression of $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ in Lemma 1, Lemma 2, and Lemma 3. This ensures that $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0$ or $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0$. This is needed because $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0$ or $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0$ is among the conditions on the sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$ in Theorem 5. This further ensures that denominator of $e(\Lambda)$ is non-zero and $e(\Lambda)$ remains defined.
Constructing an arbitrary error minimizing measurement: In Theorem 5, we saw that an arbitrary error-minimizing measurement $\Lambda$ must belong to one of $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ or $\mathcal{E}_{s}^{3}(\rho,\sigma)$. Further, the sets were written in a parameterized form in the Lemma 1, Lemma 2 and Lemma 3. Utilizing these parameters, a method to construct an arbitrary error-minimizing measurement is given in Table IV. The first column in the table states the free variables/parameters to take, and the other three give the constraints on these. The conditions for $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$ are written in second, third and fourth columns respectively. To construct an arbitrary error-minimizing measurement from $\mathcal{E}_{s}^{1}(\rho,\sigma)$, we can start from the first row, pick an arbitrary free variable satisfying the conditions stated in the second column, and proceed towards the last row in this way. After getting $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ in the last row, the error-minimizing measurement is given by $\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$. Constructing measurements from the sets $\mathcal{E}_{s}^{2}(\rho,\sigma)$ or $\mathcal{E}_{s}^{3}(\rho,\sigma)$ utilizes similar process over corresponding columns.
@@
TABLE IV: A method to construct an arbitrary error-minimizing measurement from each of the three sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$, illustrating the selection of values for various parameters
Now, we derive the acceptance for an arbitrary error minimizing measurement from the three sets in terms of the parameters given in Table IV.
Theorem 6.
For an arbitrary error minimizing measurement $\Lambda\in\mathcal{E}_{s}(\rho,\sigma)$, expressions of acceptance for the states $\rho$ and $\sigma$ are given as follows.
@@
Here, $c_{1},c_{2},c_{3}$ and $c_{r}$ are parameters as given in Table IV.
@@
$\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\rho})$ $\Rightarrow\mathrm{Tr}\left(\Gamma\rho\right)=0,$
$\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma})$ $\Rightarrow\mathrm{Tr}\left(\Gamma\sigma\right)=0,$
$\text{and }\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma})$ $\Rightarrow\mathrm{Tr}\left(\Gamma\rho\right)=0\text{ and }\mathrm{Tr}\left(\Gamma\sigma\right)=0.$
Using these, for measurements in the three sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$, and $\mathcal{E}_{s}^{3}(\rho,\sigma)$, acceptance is calculated as given below one by one.
@@
Now, we maximize the acceptance over the set $\mathcal{E}\left(\rho,\sigma,p\right)$. To do so, we maximize the expression obtained in Theorem 6 with respect to the parameters. The maximum acceptance expression is stated in the following theorem.
Theorem 7 (Maximum acceptance).
Given a pair of states as $\rho$ and $\sigma$ such that $\Pi_{\rho}\neq\Pi_{\sigma}$, then the maximum value of acceptance for states $\rho$ and $\sigma$ are summarized below in the Table V.
@@
$\Pi_{\rho}>\Pi_{\sigma}$ $1-\mathrm{Tr}\left(\Pi_{\sigma}\rho\right)$ 0
$\Pi_{\rho}<\Pi_{\sigma}$ 0 $1-\mathrm{Tr}\left(\Pi_{\rho}\sigma\right)$
@@
TABLE V: Table summarizing the maximum achievable acceptance for various cases considered in Section IV
@@
Recall from Corollary 3 that when $\Pi_{\rho}>\Pi_{\sigma}$, $\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{1}_{s}(\rho,\sigma)$. Similarly, from Corollary 4, we know that when $\Pi_{\rho}<\Pi_{\sigma}$, $\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{2}_{s}(\rho,\sigma)$. We begin with the third case, i.e., $\Pi_{\rho}\not<\Pi_{\sigma}$ and $\Pi_{\rho}\not>\Pi_{\sigma}$. In this case $\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{1}_{s}(\rho,\sigma)\cup\mathcal{E}^{2}_{s}(\rho,\sigma)\cup\mathcal{E}^{3}_{s}(\rho,\sigma)$. Starting from maximizing acceptance for the state $\rho$ over the set $\mathcal{E}_{s}(\rho,\sigma,p)$, which can be obtained by maximizing over all the three sets followed by taking the maximum of the three, we get
@@
Note that $A_{\rho}(\Lambda)=0\ \forall\ \Lambda\in\mathcal{E}_{s}^{2}(\rho,\sigma)$. Hence, we have
@@
Using the expression of acceptance from Theorem 6, we get
@@
Applying bounds on $c_{1}$ and $c_{3}$ for measurements in $\mathcal{E}_{s}^{1}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$ from Lemma 1 and Lemma 3 respectively (also given in Table IV), we get
@@
Note that $\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\leq\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{c_{r}\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\Rightarrow\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}\leq\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{c_{r}\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}$. So, we get
$\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{c_{r}\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}^{-1}\leq\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}^{-1}$
$\Rightarrow\max_{\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),\psi_{\sigma}\in\mathcal{S}(\mathrm{I}-\Pi_{\rho})}\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{c_{r}\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}^{-1}\leq\max_{\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma})}\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}^{-1}.$
Substituting above in (33), we get
@@
The equality can be proved similar to the proof of (53) in proof of Lemma C.1 of Appendix C. Note that among the all $\tilde{\psi}_{\rho}\in\mathcal{P}(\Pi_{\rho+\sigma}-\Pi_{\sigma})$, such that $\|\tilde{\psi}_{\rho}\|_{\infty}=1$, the one that maximizes $\mathrm{Tr}\left(\tilde{\psi}_{\rho}\rho\right)$ is $\tilde{\psi}_{\rho}=(\Pi_{\rho+\sigma}-\Pi_{\sigma})$. Hence, we obtain the maximum at $\psi_{\rho}=\frac{\Pi_{\rho+\sigma}-\Pi_{\sigma}}{\mathrm{Tr}\left(\Pi_{\rho+\sigma}-\Pi_{\sigma}\right)}$ and is given by
@@
The derivation of $A_{\sigma}^{s}=1-\mathrm{Tr}\left(\Pi_{\rho}\sigma\right)$ is similar, and can be obtained following the same process.
For the case $\Pi_{\rho}<\Pi_{\sigma}$ or $\Pi_{\rho}>\Pi_{\sigma}$, the maximum value of acceptance can be obtained in a similar way by maximizing over the sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$ or $\mathcal{E}_{s}^{2}(\rho,\sigma)$ respectively. ∎
@@
Note that for any desired value of acceptance, error-minimizing measurements can be designed by choosing the appropriate value of constant $c_{1},c_{2}$, or $(c_{3},c_{r})$, provided that acceptance remains less than the maximum obtained in Theorem 7.
Remark 15.
If $\Pi_{\rho}>\Pi_{\sigma}$, a measurement that maximize acceptance is $\{\Pi_{\rho}>\Pi_{\sigma},0,\mathrm{I}-\Pi_{\rho}+\Pi_{\sigma}\}$. If $\Pi_{\rho}<\Pi_{\sigma}$, a measurement that maximize acceptance is $\{0,\Pi_{\sigma}-\Pi_{\rho},\mathrm{I}+\Pi_{\rho}-\Pi_{\sigma}\}$. In the case when $\Pi_{\rho}\not<\Pi_{\sigma}$, $\Pi_{\rho}\not>\Pi_{\sigma}$, and $\Pi_{\rho}\not=\Pi_{\sigma}$, a measurement that maximize acceptance is given by $\{\Pi_{\rho+\sigma}-\Pi_{\sigma},\Pi_{\rho+\sigma}-\Pi_{\rho},\mathrm{I}+\Pi_{\rho}+\Pi_{\sigma}-2\Pi_{\rho+\sigma}\}$.
Remark 16.
For $n=2$ case, the condition for the existence of USD in Theorem 2 in [13] is given as $\textrm{supp}(\rho+\sigma)\neq\textrm{supp}(\rho)$ and $\textrm{supp}(\rho+\sigma)\neq\textrm{supp}(\sigma)$. It simplifies to the condition $\Pi_{\rho}\not=\Pi_{\sigma}$, $\Pi_{\rho}\not<\Pi_{\sigma}$, and $\Pi_{\rho}\not>\Pi_{\sigma}$. From (9), we know that zero error (i.e., USD) can be obtained for any pair of states satisfying $\Pi_{\rho}\neq\Pi_{\sigma}$. So, for a pair of states satisfying $\Pi_{\rho}<\Pi_{\sigma}$, and $\Pi_{\rho}>\Pi_{\sigma}$, USD is possible, but the maximum acceptance for one of the states vanishes as obtained in Theorem 7. So, the condition obtained in Theorem 2 in [13] is sufficient but not necessary for USD. However, the condition obtained in Theorem 2 in [13] is necessary and sufficient for USD with non-zero acceptance for both states. It is not mentioned in [13], but assumed in the proof of Theorem 2 in [13], and it becomes explicit from the maximum acceptance obtained in Theorem 7 in our paper.
V Conclusions